OpenStudy (anonymous):

Which of the following is the solution set of -2|x| < -8 {x | -4 > x > 4} {x | -4 < x < 4} {x | x < -4 or x > 4}

OpenStudy (compassionate):

Moo. I'm a cow. Do you know the first step?

OpenStudy (anonymous):

meow ima cat um no not really

OpenStudy (anonymous):

???????????????

OpenStudy (anonymous):

divide by -2 enequality sign changes \[\left| x \right|>\frac{ -8 }{-2 },or \left| x \right|>4\] \[\left| x \right|>a means x <-a and x > a\]

OpenStudy (anonymous):

use or in place of and

OpenStudy (jdoe0001):

\(|\color{red}{whatever}|< \color{blue}{something} \implies -\color{blue}{something} < \color{red}{whatever} < +\color{blue}{something}\)